Logbook for CALO
This is the logbook for the details of activities on the CALO.
Please also mention a summary in
the general logbook.
Please record activities and remarks. Details are welcomed. You may also attach pictures which is more informative.
PD readout general information
DAQ PC for PD
Cable IP: 192.168.1.111
user: calocube
Folder with binary files: /home/calocube/SPS2021/acquired_files
Folder with root files: /home/calocube/SPS2021/analysed_files
Folder with main acq and analysis code: /home/calocube/SPS2021/CaloCubeBTsoftware
Folder with main DAQ server: ~/DAQServer/build/src/
RPI4 board used as a ssh bridge for PD monitors
Wireless IP: 194.12.182.37
Cable IP: 192.168.1.107
user: pi
NAS with backup data
ssh
root@naszyxelNOSPAMPLEASE.dyndns.cern.ch
folder: /i-data/a5ceeb7e/nfs/data/TB_SPS_2021/
nfs mount:
- on DAQ PC: mount /mnt/nas
- on other PCs: sudo mount -t nfs naszyxel.dyndns.cern.ch:data/ mountdir/
Data acquisition
DAQ server for common acquisition:
~/DAQServer/build/src/calopddaqserver runCommands.txt
The standard DAQ procedure::
/home/calocube/SPS2021/CaloCubeBTsoftware/bin/acquire_TROC -C -R 2:
/home/calocube/SPS2021/CaloCubeBTsoftware/bin/acquire_TROC -R 2 -T -o /home/calocube/SPS2021/acquired_files/
Standard analysis procedure:
/home/calocube/SPS2021/CaloCubeBTsoftware/bin/analyze_alone -O /home/calocube/SPS2021/acquired_files/FILE_NAME.dat:
Standard performance monitor::
TO BE DONE
Oct 6
Oct 7
Oct 8
PD: finished mounting and testing readout. Short circuits problem found and solved. (need more details..)
Oct 9
PD: channel noise tested with the external box closed. The mean noise is < 20 ADC, while 4 channels have large noise (> 50 ADC).
The file is "20211009_154027.dat" on the RPI4 SD, folder "~/HIDRA_2.0/acquired_files/"
20:00. Acquire data using the PD system with ground muons.
Calibration file: 20211009_195607
Two large scintillators are used to select muons on the front of the calorimeter. Some calibration are events are also acquired @ 0.3 Hz.
File: 20211009_200436 20211010_032616
Oct 10
PD:
20:35. Acquire data using the PD system with ground muons.
Calibration file: 20211010_202948
Two large scintillators are used to select muons on the front of the calorimeter. Some calibration are events are also acquired @ 0.3 Hz.
File: 20211010_203553 20211011_040048
Oct 11
PD:
15:57: two calibration acquisition with PD system and
I2C trigger provided by IHEP trigger board.
file: 20211011_153110 20211011_154505
The mean noise of PD channels is < 20 ADC as expected
17:30 Acquisition with muons started, we acquired a single file: 20211011_172919.dat
Oct 14
PD:
23:45 calibration file before closing the area: 20211014_232248.root 20211014_232435.root
File with muons inside the beam area, large scintillator (10x10) is used: 20211014_232755.root
First electron @ 250
GeV file: 20211015_000302
First proton @ 350
GeV file: 20211015_001137
Starting long proton run @ 350
GeV: 20211015_011416 20211015_025359 20211015_043348 20211015_061158
Oct 15
10:30 calibration run: 20211015_103644
13:10 starting the common DAQ with protons @ 350
GeV
14:20 first acquisition with all detectors (no FIT) receiving trigger: CALOPD_RUN00035_BEAM_20211015_122209
15:30 first calibration with all detectors connected: CALOPD_RUN00037_CAL_20211015_132833
15:40 with the common DAQ for the name of the run see the table
Oct 16
Oct 17
CALOPD_RUN00116_BEAM_20211017_060424 we see some channels with adc_cn pedestal value of 18000 (suspected bad pedestal subtraction)
Oct 18
First calibration procedure done, 8 points from 0 to 3 V
free gain folder: calibration/20211018_000904
forced gain: ./calibration//20211018_002121
Oct 19
CALOPD_RUN00159_BEAM_20211018_092837.root correlation adc_cn[12][7]:adc_cn[13][7] very strange
Oct 20
Comments
--
Nicola Mori - 2021-10-15